# Network Structure Trumps Neuron Dynamics: Insights from Drosophila Connectome Simulations

### v3

@@ B.2 Model
B.2.1 Symbol Systems
The original network presentation of neural systems often uses nodes to represent the neurons and using edges to represent the links (chemical synapses, electrical synapses, etc) between neurons[4]. However, this approach has a number of limitations: first, there are usually more than one link (synapse) between two neurons[49] and the links are spatially and temporally asymmetric. To illustrate, two neurons shown in Fig.S1 have multiple synapses between them, and the dynamical properties of these different synapses may vary. Additionally, due to their spatial locations, the signal transmission delay and the position of action on the downstream neurons are also different, making it an oversimplification if the synapses are merged into one edge.
@@ B.2.1 Symbol Systems
B.2.2 Large Scale Network Communication Model
By modeling the Drosophila connectome as graph model, we introduce the large scale network communication model on this graph. We use a linear threshold activation function on each neuron. Thus, general transmission function is defined in the following manner,
@@ B.2.2 Large Scale Network Communication Model
In this paper we use threshold activation function and sigmoid function, which are two .
@@
B.2.3 Neuron Dynamic Model
@@ B.3 Evaluation Criteria
We define the consistency indicator $C_{r}$ in region $r$ which has a certain function (visual, olfactory, etc) is defined in the following manner.
$C_{r}=N_{r}^{act}/N_{r},$ (26)
where $N_{r}^{act}$ represents the number of neurons activated in area $r$ at the final state, $N_{r}$ represent the number of neurons in area $r$ and $N$ represent the total number of neurons.
Appendix C Introduction of Related Neurons
@@ Appendix D Bilateral Response under Unilateral Stimulus
In real life, the left and right hemispheres cooperate to perform complex tasks. We are interested in whether network models can simulate this cooperative effect to some extent. So we administer a unilateral stimulus to the input neurons and monitor the varying activation ratios across both hemispheres. More precisely, we independently simulate visual input neurons in the left and right hemispheres and record the rate of activated neurons in each hemisphere over time. Fig.S3 illustrates the curve depicting how the response ratio evolves with each time step. The horizontal line represents the proportion of activated neurons in all hemisphere. It is noteworthy that when simulating one hemisphere alone, visual processing neurons in both hemispheres are activated. At time step 5, the unstimulated hemisphere begins to have a activate pattern, which corresponds to the hypothesis of shallow connection[53]. Furthermore, our observations reveal a concurrent effect: the activation pattern in one hemisphere mirrors that in the unstimulated hemisphere.
These findings are reflective of reality. Although the left and right hemispheres of the brain possess certain functional specializations, they also demonstrate a synergistic effect. Typically, they collaborate to fulfill specific functions[56]. For instance, during complex cognitive tasks, both hemispheres coordinate their efforts to contribute to task completion. Even when simulating input to neurons on just one side of the brain, our model reveals similar synergistic effects, indirectly attesting to the model’s effectiveness. Besides, the mirror effect align with real situation, underscoring the brain’s resilience as a system. In instances of disease or physical damage leading to impairment in one hemisphere, this interconnected structure supports the continuation of critical functions, such as vision.
@@
Figure S3: Experiment of activation rate under single stimulus. We set threshold $\sigma_{0}=0.8$.
@@ Appendix E Impact of Perturbation Rate on Activation Areas
The main text demonstrates that with a 10% perturbation in the network, the original realistic activation patterns vanish regardless of the model used. In this section, we discuss the impact of different reconnection rates on activation patterns using the degree-preserving reconnection method.
The experiments presented in this section involve calculating patterns induced by two types of stimuli across three models under network perturbations of 1%, 3%, 5%, 10%, and 20% reconnect rates. We normalized each data point using the difference between the activation ratio of region $r$ and the average activation ratio $\bar{C}$ of the entire brain, that is to say, $\Delta C=C_{r}-\bar{C}$.
The experimental results are shown in Fig.S4, where 0 represents the original, unperturbed experimental results. It can be observed that at the 0 point, there is a significant difference in the relative activation ratios between visual and olfactory areas under specific stimulus. However, this significant realistic activation patterns which we discussed in the main text disappear with a 1% network perturbation, replaced by a roughly equal activation ratio across all regions. This indicates that even a 1% change in network structure can lead to a rapid loss of network functionality specificity. Additionally, it is noted that small-scale network perturbations (1%-5%) have a lesser impact on olfactory stimulus compared to visual stimulus. This is because the number of neurons associated with olfactory functions is significantly fewer than those related to visual functions, making it less likely for small-scale network structure changes to affect these neurons, thereby leading to a smaller impact on functionality.
At the same time, various horizontal comparisons indicate that while there are no significant differences among different models in forming activation patterns, more complex neuron models have a stronger ability to maintain their original activation patterns under network disturbances. This is because the signals generated by complex neuron models have greater redundancy and functional diversity. For instance, the LIF (Leaky Integrate-and-Fire) model has a voltage decay mechanism, where the voltage gradually decays in the absence of a signal. Even if some input signals are erroneous, as long as the proportion of errors is not very large, the decay mechanism can filter out the impact of these erroneous signals.
These results indicate that the formation of realistic activation patterns is highly sensitive to changes in network structure, with even minor perturbations capable of disrupting this network configuration. This indirectly substantiates that the authentic network structure is crucial for the formation of genuine activation patterns. Additionally, it highlights that complex functions (such as vision) are more susceptible to network disturbances.
@@ Appendix F Large Spatial Network Visualization
The visualization software of experiment observation on large spatial networks is important for researchers[57][58]. However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment. To observe directly the communication progress on this super large Drosophila connectome, we develop a large spatial network visualization software. This software will be made available to other researchers with the author’s permission.
@@ F.3 Function Display
Fig.SS5 is an example of the 3D visualization system.
@@
Figure S5: Visualization of Spatial Drosophila Brain Network
All nodes’ colors can be adjusted by visible selection or input RGB code (Fig.SS6).
Figure S6: Display of color selection
Besides, from square to round, the shape of nodes can be adjusted even real neuron (Fig.SS7).
Figure S7: Display of color selection
It also can present different single areas (Fig.SS8).
Figure S8: Display of color selection

### v4

@@ A.2 Model
A.2.1 Symbol Systems
The original network representation of neural systems often uses nodes to represent neurons and edges to represent the links (chemical synapses, electrical synapses, etc.) between neurons[3]. However, this approach has several limitations: First, multiple synapses typically exist between two neurons[17], and these synapses are spatially and temporally asymmetric. For instance, two neurons, as illustrated in Fig. S1, can have multiple synapses with varying dynamic properties. Additionally, due to differing spatial locations, signal transmission delays and the positions of action on downstream neurons also vary. Merging these synapses into a single edge oversimplifies the complexity of neural communication.
@@ A.2.1 Symbol Systems
A.2.2 Large Scale Network Communication Model (LSNC)
By modeling the Drosophila connectome as graph, we introduce the large scale network communication model on this graph. The general transmission function is defined in the following manner,
@@ A.2.2 Large Scale Network Communication Model (LSNC)
In this paper we use threshold activation function and sigmoid function for comparison.
@@
A.2.3 Neuron Dynamic Model
@@ A.3 Evaluation Criteria
We define the consistency indicator $C$ in a area which has a certain function (vision, olfaction, etc) is defined in the following manner.
$C=N^{act}/N,$ (14)
where $N^{act}$ represents the number of neurons activated in this area at the final state, $N$ represent the number of neurons in this area and $N$ represent the total number of neurons.
Appendix B Introduction of Related Neurons
@@
Tab. 3 represents the numbers of specific neurons of each hemisphere in related areas. It presents that the number of neurons in specific areas of the Drosophila brain connectome is mirror-symmetrical. Additionally, visual neurons constitute the largest proportion among the connectome’s neurons (approximately one-twice). Neurons in the olfactory-related regions, though comprising a smaller proportion (for instance, the ALON region has only 8 neurons on each side), play a critical role in olfactory perception.
Table 3: Node numbers in related areas Drosophila Brain Connectcome Graph
total (left) total (right) optic (left) optic (right)
$60834$ $63144$ $36131$ $37524$
visual projection (left) visual projection (right) ALIN (left) ALIN (right)
$3117$ $3117$ $12$ $12$
ALLN (left) ALLN (right) ALON (left) ALON (right)
$196$ $194$ $8$ $8$
ALPN (left) ALPN (right)
$310$ $314$
@@
Appendix C Experimental Settings
The basic methodology of the experiment is as follows. We apply a continuous stimulus signal (binary 0-1 signal) to a subset of input neurons (i.e., visual or olfactory neurons) in the brain. We then compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons. We give 2 types of input neurons (visual, olfactory) stimulus and record the activation patterns of 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory). When we give neuron a stimulus, we set the state of neuron as 1. The evaluation criterion involves calculating the ratio of neurons that should be activated to those that are actually activated in a specific area (see Equ. 14 in Appendix A).
Given that the brain functions similarly to a shallow neural network[22], we document outcomes at an iteration step size of 5. Since the input is constant and no randomization parameters are introduced, the results of each experiment are consistent, eliminating the need for repetition. For detailed information on related neurons, please read Appendix B.
In the linear threshold model, we set $\sigma_{0}=0.8$. For the simulation of the LIF model, numerous hyperparameters are required. All the hyperparameters used are listed in the following table. Biological hyperparameters are set according to real biological experimental tests[23]. The current proportion constant is the optimal parameter obtained through equidistant search.
Table 4: List of Hyperparameters
Hyperparameter Description Value
$\tau$ Membrane time constant 15.0 ms
$V_{\text{rest}}$ Resting membrane potential -65 mV
$V_{\text{th}}$ Spike threshold -45 mV
[…]
$dt$ time step 0.55 ms
$\alpha$ Current proportion const 0.03/0.045 (visual/olfactory)
To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks. Specifically, we perform degree-preserving rewiring on a certain proportion of edges in the brain, which means perturbing the network structure while ensuring that the degree of each node remains unchanged. In Fig. 2 we set reconnect rate $p=20\%$. For its sensitivity analysis, please read Appendix E.
@@ Appendix D Bilateral Response under Unilateral Stimulus
In real life, the left and right hemispheres cooperate to perform complex tasks. To examine whether network models can simulate this cooperation, we administered a unilateral stimulus to input neurons and monitored the varying activation ratios across both hemispheres. Specifically, we independently simulated visual input neurons in the left and right hemispheres and recorded the rate of activated neurons in each hemisphere over time. Second, we investigate whether local attack to one side of the network—analogous to real-world conditions such as tumors and other diseases—affects this bilateral response process, and how the extent of such damage influences the synergistic response.
The specific methods of inducing attack are as follows: We first identify two symmetrical coordinates on both hemispheres of the brain. Using these coordinates as the center, we remove nodes within varying radii and subsequently record the response curves for each hemisphere-related region.
Fig. S3 illustrates the specific length measurements of the Drosophila connectome and the regions examined in this study. The purple and blue cells represent neurons associated with visual processing. Spheres $S1$, $S2$, $S3$, and $S4$ each represent four spherical regions subjected to localized attacks. Sphere $S1$ and $S2$ encompass approximately 2,000 neurons each, while sphere $S3$ and $S4$ each cover approximately 12,000 neurons.
(a) Geometric dimensions of Drosophila connectome
(b) Locations of targeted regions
Figure S3: Schematic diagram of the specific length measurements of the Drosophila connectome and the location of regions deleted. The units of center and r in labels are micrometers.
Fig. S4 illustrates the curve depicting how the response ratio evolves with each time step. We use $C_{avg}$ to represent the average activated rate of areas related to vision. Notably, when stimulating one hemisphere alone, visual processing neurons in both hemispheres are activated (see curve regular left and right in Fig. S4(a) and (b)).
[…]
Below, we further explain the reasons for the significant delay in responses to unilateral large-scale regional attacks on this side’s stimuli (e.g., in Fig. S4(a), where the local damaged area is the $S2$ region of the left hemisphere, and stimuli on the left side cause responses in both the left and right hemispheres (light purple and dark purple curves) to be delayed). Calculations show that there are 19,031 synapses from the left hemisphere’s visual input area to the left hemisphere’s visual signal processing area (optic, visual_projection), while there are only 18 synapses from the left hemisphere’s visual input area to the right hemisphere’s visual signal processing area. Furthermore, there are 338,456 synapses from the left hemisphere’s visual processing area to the right hemisphere’s visual signal processing area. This implies that, primarily, the unilateral visual input area essentially does not interact with the contralateral visual signal processing area, and interactions occur indirectly through the ipsilateral visual processing area. Consequently, when the visual signal processing area of this side is damaged, the reduction in the number of neurons makes it challenging for the remaining neurons to reach an excitatory state, causing delays in the ipsilateral visual signal processing area. This delay, in turn, leads to delayed responses in the contralateral visual signal processing area.
@@
Figure S4: Experiment of activation rate under single stimulus. All responses on the left hemisphere are marked with light-colored lines, and all responses on the right hemisphere are marked with dark-colored lines. We set threshold $\sigma_{0}=0.95$.
@@ Appendix E Impact of Perturbation Rate on Activation Areas
In this section, we discuss the impact of different reconnection rates on activation patterns using the degree-preserving reconnection method.
The experiments presented in this section involve calculating patterns induced by two types of stimuli across three models under network perturbations of 1‰, 5‰, 1%, 3%, 5%, 10%, and 20% reconnect rates. Let $C$ be the average activation ratio of an area. We normalized each data point using the difference between the activation ratio of this area and the average activation ratio $\bar{C}$ of the entire brain, that is to say, $\Delta C=C-\bar{C}$.
The experimental results are shown in Fig. S5, where 0 represents the original, unperturbed experimental results. At the 0 point, there is a significant difference in the relative activation ratios between visual and olfactory areas under specific stimuli. However, this significant activation pattern, discussed in the main text, disappears with a 1‰’s network perturbation, replaced by a roughly equal activation ratio across all regions. This indicates that even a 1‰’s change in network structure can lead to a rapid loss of network functionality specificity. Additionally, small-scale network perturbations (1‰-5%) have a lesser impact on olfactory stimuli compared to visual stimuli. This is because the number of neurons associated with olfactory functions is significantly fewer than those related to visual functions, making it less likely for small-scale network structure changes to affect these neurons, thereby leading to a smaller impact on functionality.
Horizontal comparisons indicate that while there are no significant differences among different models in forming activation patterns, more complex neuron models better maintain their original activation patterns under network disturbances. This is because signals generated by complex neuron models have greater redundancy and functional diversity. For instance, the LIF (Leaky Integrate-and-Fire) model has a voltage decay mechanism, where the voltage gradually decays in the absence of a signal. Even if some input signals are erroneous, as long as the proportion of errors is not very large, the decay mechanism can filter out the impact of these erroneous signals.
@@ Appendix F Large Spatial Network Visualization
The visualization software of experiment observation on large spatial networks is important for researchers. However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment. To observe directly the communication progress on this super large Drosophila connectome, we develop a large spatial network visualization software. This software will be made available to other researchers with the author’s permission.
@@ F.3 Function Display
Fig. S6 is an example of the 3D visualization system.
@@
Figure S6: Visualization of spatial Drosophila brain network.
All nodes’ colors can be adjusted by visible selection or input RGB code (Fig. S7).
Figure S7: Display of color selection.
Besides, from square to round, the shape of nodes can be adjusted even real neuron (Fig. S8).
Figure S8: Display of nodes size adjusting.
It also can present different single areas (Fig. S9).
Figure S9: Display of area selection.